You have a planning horizon of 8 years and wish to immunize your investment over that horizon. To do so, you have two financial instruments available: A and B. Instrument A has a duration of 2 years, and instrument B has a duration of 13 years. Find the proportion of your money that you need to invest in instrument A in order to achieve your immunization goal.
Answer #1

Let Proportion of Money Invested in instrument A=x
Proportion of Money Invested in instrument B=1-x
Duration of portfolio=x*2+(1-x)*13=8
2x+13-13x =8
11x =13-8 =5
x =5/11
Proportion invested in A =5/11 =0.45
Proportion invested in B =1-x =1-5/11 =6/11=0.55